OpenStudy (anonymous):

Find the zeros of the function. f(x) = 2x3 - 12x2 + 16x

OpenStudy (anonymous):

First factor out 2x 2x(x^2-6x+8) factor inside the parentheses 2x(x-4)(x-2) so 2x=0 and x-4=0 and x-2=0

OpenStudy (anonymous):

thank you! (: